A sample of vinaigrette salad dressing has a total mass of 188 g

188 g , a total volume of 196 cm

3

196⁢ cm3 , and consists only of oil (density=0.882 g/cm

3

)

(density=0.882⁢ g/cm3) and vinegar (density=1.056 g/cm

3

)

(density=1.056⁢ g/cm3) .

What is the volume of each component in this mixture? Assume that the volumes are additive.


1
Expert's answer
2021-09-03T01:47:07-0400

"m=188g, v=196cm^3,density_1=0.882g\/cm^3,density_2=1.056g\/cm^3"

"v_1=?, v_2=?"

"v=m_1\/density_1+m_2\/density_2=196"

"m=m_1+m_2=188, m_1=188-m_2"

"m_1\/0.882+m_2\/1.056=196"

"1.056m_1+0.882m_2=196\times 0.882\times 1.056"

"1.056(188-m_2)+0.882m_2=182.553"

"198.528-1.056m_2+0.882m_2=182.553"

"15.975=0.174m_2"

"m_2=91.81g"

"v_2=m_2\/density_2=91.81\/1.056=86.94cm^3"

"v_1=196-86.94=109.06cm^3"

volume of oil="109.06cm^3"

volume of vinegar="86.94cm^3"